OpenStudy (anonymous):

Use the binomial probability formula to fithe probability of x success given the probability p of success on a single trial. n=30; x=5; p=1/5

OpenStudy (campbell_st):

so the formula is \[P(E) = ^nC_{r} p^rq^{n - r}\] so in your question q = 4/5 C is for combinations so \[P(5) = ^{30}C_{5} (\frac{1}{5})^5 \times (\frac{4}{5})^{30-5}\] just calculate it out.